if y<-7, evaluate the expression
(absolute value of y+7) over y+7

Answer by math problem solving(30)

Note the fact that:
|a|= - a, if a < 0
As y< - 7, so y+7 < 0
Thus |y+7| = - (y+7)
Therefore:
|y+7|/(y+7) = - (y+7)/(y+7) = - 1